Finally - Attachment #1. This plot uses 16 Hz EPICS data. All y-axes are uncalibrated for now, but TRX/TRY are normalized such that the POX/POY lock yields a transmission of 1. CARM UGF is only ~3 kHz, no boosts were turned on yet.
Attachment #2 and Attachment #3 are phone photos of the camera images of the various ports. After some alignment work, the transmitted arm powers were ~200, i.e. PRG ~10. fwiw, this is the darkest i've ever seen the 40m dark port. c.f. 2016. Of course, the exposure time / ND filter / light levels could all have changed.
This work was possible during the daytime (~6pm PDT), but probably only because it was Sunday. The other rate limiting factor here is the franky terrible IMC duty cycle. TBH, I didn't honestly expect to get so far and ran out of time, but I think the next steps are:
As usual, I would like to request that we don't change the IFO as far as possible until the BHD vent, i found it pretty difficult to get here.
Attachment #4 now shows the measured DARM OLTF when DARM is entirely on AS55_Q control. UGF is ~120 Hz and the phase margin is ~30 deg, seems okay for a first attempt. I'll now need to infer the OLTF over a wider range of frequencies by lining this measurement up with some model, so that I can undo the loop in plotting the DARM ASD.